OpenStudy (selena2345):

You are playing a board game with a friend.You draw one card each turn. If you draw a positive integer you move forward. If You draw a negative integer you move backward. You draw five cards shown on your first five turns. How far and in what direction have you moved along the board after these five turns?

OpenStudy (anonymous):

what are the five cards

OpenStudy (selena2345):

31 5 -4 -3

OpenStudy (selena2345):

3 1 5 -4 -3

OpenStudy (anonymous):

so your 5 numbers are3,1,5,-4,-3??? is that correct

OpenStudy (selena2345):

Yes that is correct

OpenStudy (anonymous):

ok easy we start going forward 3+1=4 ok now forward again 4+5=9 now back four squares 9-4=5 now back three more squares 5-3=2 you have moved two steps forward on your turn congratulations!

OpenStudy (selena2345):

Thank You countonme123

OpenStudy (anonymous):

sure can i have a medal

OpenStudy (anonymous):

thanks if you ever need anything just ask me you are my fan :)